# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ 4 Simulation Studies
In Simulation 1, we explore how positivity violations exacerbate bias in causal estimates, particularly when the outcome models are misspecified. Simulation 2 examines the impact of positivity violations on causal estimation by comparing scenarios with and without positivity violations. Simulation 3 investigates the methods in a more complex setting, incorporating multiple confounders and intricate data-generating processes.
@@ 4.1.1 Simulation 1: Correct model specification versus misspecification under positivity violations
In Simulation 1A, the outcome function is linear. In Simulation 1B, the outcome function is sinusoidal, and in Simulation 1C, it is logarithmic. These two non-linear outcome functions in Simulations 1B and 1C pose a greater challenge for estimation, particularly in regions where data is sparse, i.e. suffering from positivity violations, leading to potential estimation difficulties. In contrast, the linear outcome function in Simulation 1A is expected to be easier to estimate, even in the presence of data sparsity, as the extrapolation will work perfectly.
@@ 4.2 Estimation and Evaluation
We set the sample size to $n=1000$ for all simulation studies. The true value of each estimand $m^{\varepsilon}$, where $\varepsilon\in\{\text{standard},\text{feasible},\text{trimming}\}$, is computed by intervening on the data-generating process using the specified actions, with a large sample size of N=100, [^0] N=100{,}000. For estimation (see Algorithm 1 for details), we apply the parametric g-computation plug-in estimator in Simulations 1 and 2, and the HAL-based estimator in Simulation 3. The method for conditional density estimation varies by scenario: parametric linear regression (Simulation 1), nonparametric kernel methods (Simulation 2, using the R package np [12]), and Super Learner (Simulation 3; see Appendix C). We assess bias over $R=1000$ Monte Carlo replications, defining absolute bias at level $a$ as:
$\text{Absolute Bias}(a)=\dfrac{1}{R}\sum_{r=1}^{R}|\hat{m}^{\varepsilon}_{r}(a)-m^{\varepsilon}(a)|.$
with $m^{\varepsilon}$ the true target parameter.
@@ 4.3.1 Simulation 1
Figure 3c examines the impact of non-overlap and model misspecification on the estimation of three estimands under linear, logarithmic, and sinusoidal outcome functions. In regions where the non-overlap ratio is near zero, counterfactual estimates for the three estimands are nearly identical across all outcome functions, as intended by the simulation design. At the extremes, however, the estimands diverge. The CDRC with the most feasible intervention establishes practical thresholds, avoiding unrealistic extrapolation into sparsely supported regions. The trimming estimand, by contrast, restricts the population to those for whom the intervention is feasible.
@@
The most feasible intervention consistently achieves lower absolute bias, particularly when outcome functions are misspecified and extrapolation is unreliable. By focusing on well-supported regions, it avoids large biases and maintains robust estimates even in areas with low non-overlap. Conversely, the trimming estimand exhibits high variability and substantial absolute bias, even in the linear case, due to its restrictive population definition.
@@ 4.3.2 Simulation 2
Figure 4c shows the true estimand values for both the standard CDRC, the CDRC with the most feasible intervention and trimming estimand. In Simulation 2A, where non-overlap is absent, the three true curves are nearly identical because the intervention is feasible for all individuals. However, in Simulation 2B, the estimands diverge outside the central region, as the most feasible intervention shifts toward the center due to limited feasibility at extreme values, and the trimming estimand exhibits more radical counterfactual shifts. For all three curves, the estimates are very close to the true values, as the true outcome function is designed to be straightforward.
Figure 4d compares the absolute bias between Simulations 2A and 2B. Compared to Simulation 2A, Simulation 2B exhibits positivity violation issues, leading to higher absolute bias for all estimands, particularly in regions with high non-overlap. Nonetheless, the most feasible intervention shows lower bias under both conditions compared to the standard estimand and does not exceed the bias level of the standard estimand when positivity violations are minimal, as observed in Simulation 2A. The trimming estimand performs well at times, but it is unstable in certain areas.
@@ 4.3.3 Simulation 3
Figure 5b presents the true and estimated values for the three estimands. In regions where the non-overlap ratio is zero, the three estimands are indistinguishable. However, the standard estimand increases dramatically near an intervention value of 0. As indicated by the non-overlap ratio, such values are nearly impossible to intervene on. In the data-generating process (see Appendix B.3), the truncated normal distribution excludes values below $0.2032$, causing the true curve of the trimmed estimand to terminate near this intervention value. This indicates that intervention values below $0.2032$ are strictly infeasible, representing a theoretical positivity violations and thus making counterfactual curve of trimming estimand unidentifiable below that threshold. Unlike the standard estimand, the proposed most feasible intervention estimand does not exhibit extreme values in regions with very low intervention, making it more robust and less prone to positivity issues.
Figure 5c demonstrates that the most feasible intervention estimand consistently achieves lower absolute bias, particularly in regions with positivity violations. It effectively avoids large biases and maintains robust estimates even in areas with low or no non-overlap. While the trimming estimand also reduces bias, it becomes not estimable when the non-overlap ratio approaches $1$, as no samples are available in such scenarios.
@@ 5 Data Analysis
For illustration, we use baseline data and data from the scheduled visit at $t=36$ (weeks). The relevant part of the assumed data-generating process is, in line with previous studies [30, 18, 31], summarized by the directed acyclic graph (DAG) in Figure 6:
[…]
Figure 6: Directed acyclic graph (DAG) for CHAPAS-3 trial described in Section 5. NRTI refers to the Nucleoside Reverse Transcriptase Inhibitor drug drug component, sex, age, genotype relates to how fast the drug is cleared as determined by the 516G and 983T polymorphisms in the CYP2B6 gene [4], non-adherence describes any signs of non-adherence as measured through a medication event monitoring system [30] as well as weight. Subscripts describe the the week of measurement.
Our analysis illustrates the ideas based on a complete case analysis of all measured variables at $t=36$ (weeks) represented in the DAG $(n=83)$. The analysis was conducted using three support levels for detecting non-overlap: $99\%$, $95\%$, and $90\%$. Our target estimand is the standard causal concentration-response curve, defined in Section 3.1. We compare this estimand to the proposed estimand with most feasible intervention strategy, described in Section 3.3.3, as well as the trimming estimand in Section 3.3.4. The estimation process closely follows the detailed methodology described in Section 3.4 and algorithm outlined in Algorithm 1.
Note: To quantify uncertainty, obtaining reliable compatibility intervals is challenging in this setting, both because the functional parameter defined in Equation (16) is not pathwise differentiable and because of uncertainty around ideal model specification. Since bootstrap inference may be unreliable in such nonregular problems, we report only point estimates in the application and acknowledge this limitation.
@@ 5.1 Results
Figure 7: Results of the data analysis
Figure 7a illustrates the proposed non-overlap ratios as diagnostics for identifying positivity violations in continuous interventions. These ratios, calculated as a function of the intervention $\text{EFV}_{36}$, represent the proportion of the population where the intervention is not feasible, based on 3 support levels of non-overlap detection.
The non-overlap ratios remain low across all support levels within the central intervention range (approximately $1$ to $3.5$ mg/L), suggesting minimal violations of positivity in this region and supporting reliable causal estimation. However, at the boundaries of the intervention range—below $1$ mg/L or above $3.5$ mg/L—the ratios increase, indicating more frequent violations. Notably, near 0 mg/L, the non-overlap ratio rises sharply, signaling strong violations of the positivity assumption, consistent with the fact that concentrations this low are virtually impossible to observe.
Figure 7b show the estimated curves of the most feasible intervention across three support levels compared with the standard causal concentration-response curve. In regions with positivity violations, using the most feasible intervention strategy differs from the CDRC, while closely aligning with standard estimates in well-supported regions. This means whenever a concentration value seems unlikely or implausible, we do not enforce this intervention but rather use the closest value that is possible. This may happen, for example, if there are patients with a very slow metabolism who will never be able to clear the drug fast enough to ever achieve concentrations under $1$ mg/L.
Figure 7c compares all estimated causal curves across three support levels. As the support level increases, more positivity violations are tolerated for both the feasible intervention and triming estimands. However, the triming estimand exhibits instability in regions with strong positivity violations, suggesting that it may not be the optimal approach in this analysis.
@@ 6.1 Summary
Positivity violations are a persistent challenge in causal inference with continuous interventions, particularly when outcome regressions must be extrapolated from observed to unobserved regions. Such violations arise when certain intervention levels are insufficiently represented in the data, undermining both the identification and estimation of causal effects.
We address this challenge with a data-adaptive framework that employs an $L$-specific density threshold $f_{\alpha}$ to define highest density regions $\mathcal{A}_{\alpha}(\mathbf{l})$. Conceptually, this idea is related to Modified Treatment Policies (MTPs), since both approaches avoid assigning interventions to strata with poor support. The objectives, however, differ: MTPs define modified treatment (often based on natural values) to ensure feasibility, whereas our method preserves the causal dose–response curve (CDRC) as much as possible. As Ring et al. [28] note, even MTPs may not fully resolve positivity violations when the modified treatment lacks adequate support.
Our approach also contrasts with the weighted CDRC proposed by Schomaker et al. [31]. Both approaches address positivity violations via conditional densities, but they differ in how the estimand is adapted. First, we employ a unit-specific, data-adaptive threshold, whereas Schomaker et al. use a fixed cutoff $c$, which may be unsuitable across intervention ranges and covariate strata. Second, our framework serves dual purposes: it functions both as a diagnostic tool (see Section 2.3) and as a procedure for defining a feasible intervention rule. A further distinction lies in interpretability. Our method modifies intervention assignments explicitly (to the “closest feasible” value) and targets the clearly defined quantity $E(Y^{\mathop{}\!\mathrm{d}})$. In contrast, the weighted CDRC redefines the estimand implicitly through a projection within the expectation, thereby changing the target parameter.
@@
A key consequence of positivity violations is increased estimation bias—even when models are correctly specified—because estimation relies on extrapolation into unsupported regions. The issue is therefore not merely one of model misspecification but one of fundamental identifiability. Our data-adaptive intervention framework addresses this challenge by reassigning infeasible interventions to regions with adequate empirical support, thereby avoiding extrapolation and ensuring more reliable estimation.
@@ 6.2 Limitations
Estimating the full CDRC under continuous interventions faces several fundamental limitations. First, under mild smoothness conditions, the CDRC is non-pathwise differentiable [8]. This rules out the use of influence-function-based estimators such as TMLE or doubly robust procedures, unless additional structural assumptions are imposed. While some recent approaches achieve robustness by assuming differentiability [22], monotonicity [36], or smooth invertibility [9], such assumptions may be difficult to defend in applied contexts where the estimand is determined by the scientific question. Thus, double robustness and semiparametric efficiency are generally unattainable in this setting.
@@
Second, the data-adaptive nature of our proposed intervention introduces dependence between the observed data and the estimand. As noted by Hubbard et al. [19], this dependence violates standard independence assumptions and may compromise the validity of inference procedures such as the bootstrap. Sample splitting or cross-validation can help restore approximate independence by separating the construction of the data-adaptive intervention rule from estimation of the target parameter, but these remedies introduce practical complexity: they yield multiple, split-specific estimands and complicate interpretability.
@@
Third, the performance of our approach depends on the stability of the data-adaptive rule itself. In particular, the choice of the support level $\alpha$, the method for estimating highest density regions, and the specification of the substitution function all influence the resulting estimand. Instability in these components may introduce variability or sensitivity in finite samples, raising concerns about robustness and reproducibility across different applications.
@@
In light of these challenges, our work pursues a complementary goal: preserving the original CDRC estimand while addressing positivity violations through data-adaptive interventions. This yields a natural plug-in estimator that avoids extrapolation and provides valid inference under correct model specification, provided the data-adaptive rule is reasonably stable. Although our approach does not deliver double robustness, it enhances interpretability and resilience to misspecification by focusing estimation on well-supported regions. To evaluate the impact of data-adaptivity on inference, we conducted a simulation study (Appendix D) comparing bootstrap coverage with and without sample splitting. The results suggest that, in well-supported regions, bootstrap intervals remain reasonably accurate even without splitting.
@@
Finally, although our method provides a principled approach for substituting infeasible interventions with their most feasible counterparts—thereby preserving fidelity to the standard CDRC and enhancing interpretability—important avenues remain for future work. In particular, further research is needed to extend the framework to longitudinal settings, refine the design of substitution rules, and develop more robust techniques for estimating support regions, especially in high-dimensional and complex treatment spaces.

### v3

@@ 5 Simulation Studies
In Simulation 1, we explore how positivity violations exacerbate bias in estimation, particularly when the outcome models are misspecified. Simulation 2 examines the impact of positivity violations on estimation by comparing scenarios with and without positivity violations. Simulation 3 investigates the methods in a more complex setting, incorporating multiple confounders and intricate data-generating processes.
@@ 5.1.1 Simulation 1: Correct model specification versus misspecification under positivity violations
In Simulation 1A, the true outcome function is linear. In Simulation 1B, the true outcome function is sinusoidal, and in Simulation 1C, it is logarithmic. These two non-linear outcome functions in Simulations 1B and 1C pose a greater challenge for estimation, particularly in regions where data is sparse, i.e. suffering from positivity violations, leading to potential estimation difficulties. In contrast, the linear outcome function in Simulation 1A is expected to be easier to estimate, even in the presence of data sparsity, as the extrapolation will work perfectly.
@@ 5.2 Estimation and Evaluation
We set the sample size to $n=1000$ for all simulation studies. The true value of each estimand $m^{\dagger}_{0}$, where $\dagger\in\{\text{standard},\text{feasible},\text{trim}\}$, is approximated by intervening on the data-generating process under the corresponding action using a large Monte Carlo sample of size N=100, [^0] N=100{,}000. For estimation (see Algorithm 1 for details), we apply the parametric g-computation plug-in estimator in Simulations 1 and 2, and the HAL-based estimator in Simulation 3. The method for conditional density estimation varies by scenario: parametric linear regression (Simulation 1), nonparametric kernel methods (Simulation 2, using the R package np [12]), and hazard-binning (Simulation 3; see Appendix 8.2). We assess bias over $R=1000$ Monte Carlo replications, defining absolute bias at level $a$ as
$\text{Absolute Bias}(a)=\dfrac{1}{R}\sum_{r=1}^{R}\left|\hat{m}^{\dagger}_{r}(a)-m^{\dagger}_{0}(a)\right|.$
@@ 5.3.1 Simulation 1
Figure 3c examines the impact of non-overlap and model misspecification on the estimation of three estimands under linear, logarithmic, and sinusoidal outcome functions. In regions where the non-overlap ratio is near zero, counterfactual estimates for the three estimands are nearly identical across all outcome functions, as intended by the simulation design. At the extremes, however, the estimands diverge. The FDRC establishes practical thresholds, avoiding unrealistic extrapolation into sparsely supported regions. The trimming estimand, by contrast, restricts the population to those for whom the intervention is feasible.
@@
The FDRC achieves lower absolute bias, particularly when outcome functions are misspecified and extrapolation is unreliable. By focusing on well-supported regions, it avoids large biases and maintains robust estimates even in areas with low non-overlap. Conversely, the trimming estimand exhibits high variability and substantial absolute bias, even in the linear case, due to its restrictive population definition.
@@ 5.3.2 Simulation 2
Figure 4c shows the true estimand values for both the standard CDRC, FDRC and trimming estimand. In Simulation 2A, where non-overlap is absent, the three true curves are nearly identical because the intervention is feasible for all individuals. However, in Simulation 2B, the estimands diverge outside the central region, as the most feasible intervention shifts toward the center due to limited feasibility at extreme values, and the trimming estimand exhibits more radical counterfactual shifts. For all three curves, the estimates are very close to the true values, as the true outcome function is designed to be straightforward.
Figure 4d compares the absolute bias between Simulations 2A and 2B. Compared to Simulation 2A, Simulation 2B exhibits positivity violation issues, leading to higher absolute bias for all estimands, particularly in regions with high non-overlap. Nonetheless, the FDRC shows lower bias under both conditions compared to the standard estimand and does not exceed the bias level of the standard estimand when positivity violations are minimal, as observed in Simulation 2A. The trimming estimand performs well at times, but it is unstable in certain areas.
@@ 5.3.3 Simulation 3
Figure 5b presents the true and estimated values for the three estimands. In regions where the non-overlap ratio is zero, the three estimated curves are indistinguishable. However, the standard CDRC increases dramatically near an intervention value of 0. As indicated by the non-overlap ratio, such values are nearly impossible to intervene on. In the data-generating process (see Appendix B.3), the truncated normal distribution excludes values below $0.2032$, causing the true curve of the trimmed estimand to terminate near this intervention value. This indicates that intervention values below $0.2032$ are strictly infeasible, representing a theoretical positivity violations and thus making counterfactual curve of trimming estimand unidentifiable below that threshold. Unlike the standard CDRC, the proposed FDRC does not exhibit extreme values in regions with very low intervention, making it more robust and less prone to positivity issues.
Figure 5c demonstrates that the FDRC consistently achieves lower absolute bias, particularly in regions with positivity violations. It effectively avoids large biases and maintains robust estimates even in areas with low or no non-overlap. While the trimming estimand also reduces bias, it becomes not estimable when the non-overlap ratio approaches $1$, as no samples are available in such scenarios.
@@ 6 Data Analysis
For illustration, we use baseline data and data from the scheduled visit at $t=36$ (weeks). The relevant part of the assumed data-generating process follows previous studies [30, 18, 31].
Our analysis illustrates the ideas based on a complete case analysis of all measured variables at $t=36$ (weeks) represented in the DAG $(n=83)$. The analysis was conducted using three support levels for detecting non-overlap: $99\%$, $95\%$, and $90\%$. Our reference target estimand is the standard causal concentration-response curve, defined in Section 4.1. We compare this estimand to the proposed FDRC, described in Section 4.3, as well as the trimming estimand in Section 4.4. The estimation process closely follows the detailed methodology described in Section 4.5 and algorithm outlined in Algorithm 1.
@@ 6.1 Results
Figure 6: Results of the data analysis. In panels (b)–(c), the x-axis shows the efavirenz concentration level we aim for (the target/attempted concentration). The feasible curve answers: “If we try to set everyone to this concentration, but for some children it is not realistically attainable, what would the outcome look like after moving those children to their nearest attainable concentration?” The trimmed curve answers: “What would the outcome look like at this concentration among the children for whom this concentration appears supported by the data?” The standard CDRC is shown at the same target concentration values for comparison.
Figure 6a reports the proposed non-overlap ratio as a diagnostic for positivity violations under continuous interventions. For each target intervention value $\text{EFV}_{36}=a$, the ratio estimates the fraction of children for whom that target appears unsupported given their baseline covariates, shown for three support levels ($90\%$, $95\%$, and $99\%$).
Across the central range of concentrations (approximately $1$ to $3.5$ mg/L), the non-overlap ratio is close to zero at all support levels, indicating that these targets are broadly supported and that causal estimation is least sensitive to positivity issues. In contrast, the ratio increases at the boundaries—below $1$ mg/L and above $3.5$ mg/L—indicating that these targets are unsupported for a nontrivial subset of the population. The sharp rise near 0 mg/L is consistent with concentrations in this region being extremely rare in the data.
Figure 6b shows the estimated feasible dose–response curves at three support levels alongside the standard causal concentration–response curve (CDRC). In well-supported regions the feasible and standard curves are similar, while they diverge near the boundaries where non-overlap is detected. This behavior reflects the definition of the feasible strategy: for a given target concentration on the x-axis, children for whom that target is unsupported are mapped to the closest supported concentration before evaluating the outcome regression. In this setting, such adjustments are most relevant for extreme targets that some clinical/metabolic profiles are unlikely to attain.
Figure 6c compares all estimated curves by support level. As the support level increases, the procedure tolerates more non-overlap, so feasible and trimming curves can differ over a wider range of targets. In this analysis, the estimated trimming curve becomes unstable in regions with strong non-overlap, suggesting limited practical support at the extremes.
Overall, in this HIV application the feasible curve should be read as an attempted-concentration curve: the x-axis is the EFV36 concentration level we try to set for everyone. For each attempted value $a$, the curve answers the following question: if we aim for $a$ for every child, but $a$ is not realistically attainable for some children given their covariates (as indicated by the estimated non-overlap), what would the average viral-failure probability look like after assigning those children to their nearest attainable concentration? The resulting curve is plotted against the attempted value $a$, so differences from the standard CDRC occur primarily where the data indicate that the literal intervention “set $A=a$ for everyone” is not well supported.
@@ 7.1 Summary
Positivity violations are a persistent challenge in causal inference with continuous interventions, particularly particularly when estimation requires extrapolation from observed to unobserved regions. Such violations arise when certain intervention levels are insufficiently represented in the data, undermining both the identification and estimation of causal effects.
We address this challenge by using an $\mathbf{L}$-specific density cutoff $f_{\alpha}(\mathbf{l})$ to define conditional highest-density support sets $\mathcal{A}_{\alpha}(\mathbf{l})$. This is closely related in spirit to Modified Treatment Policies (MTPs): both adapt interventions to the observed support structure and therefore reduce reliance on extrapolation in low-overlap strata. The distinction is in the target and construction. MTPs typically define a prespecified shift/transform of treatment (anchored to observed treatment values), while our framework is indexed by a target level a and then enforces feasibility by restricting or projecting to $\mathcal{A}_{\alpha}(\mathbf{l})$, with the goal of preserving comparability to the standard CDRC as much as possible. As emphasized by Ring et al. [28], positivity concerns for MTPs can remain if the induced intervention values themselves are weakly supported; this same principle motivates our explicit non-overlap diagnostics and HDR-based restriction.
Our approach also contrasts with the weighted CDRC proposed by Schomaker et al. [31]. Both approaches address positivity violations via conditional densities, but they differ in how the estimand is adapted. First, we use a covariate-specific support set (indexed by $\alpha$) rather than a fixed cutoff $c$, whereas Schomaker et al. use a fixed cutoff $c$, which may be unsuitable across intervention ranges and covariate strata. Second, our framework serves dual purposes: it provides both overlap diagnostics (see Section 3.3) and a principled way to define a feasible intervention rule. A further distinction lies in interpretability. Our method modifies intervention assignments explicitly (to the “closest feasible” value) and targets the clearly defined quantity $E(Y^{d_{0,\alpha}(a,\cdot)})$. In contrast, the weighted CDRC redefines the estimand implicitly through a projection inside the expectation, which can make the target parameter less transparent.
@@
A key consequence of limited overlap is increased finite-sample bias and sensitivity to modeling choices, because inference effectively relies on extrapolation into poorly supported regions. Thus, the problem is not only model misspecification but limited empirical support; under strict violations, identification may fail altogether. Our intervention framework addresses this challenge by reassigning infeasible interventions to regions with adequate support, thereby reducing reliance on extrapolation and typically improving finite-sample performance.
@@ 7.2 Limitations
Estimating the full CDRC under continuous interventions faces several important limitations. First, under mild smoothness conditions, the CDRC is non-pathwise differentiable [8]. This makes it more challenging to apply standard influence-function-based estimators such as TMLE or doubly robust procedures in the fully nonparametric model, and typically motivates introducing additional structure. Several recent approaches obtain robustness by leveraging assumptions such as differentiability [22], monotonicity [37], or smooth invertibility [9]. While these assumptions can be reasonable in many applications, they may not always hold or may require careful justification depending on the scientific context. Accordingly, achieving double robustness or semiparametric efficiency often relies on working under a more restricted model.
@@
Second, our proposed feasible intervention strategy introduces dependence between the observed data and the estimand. As noted by Hubbard et al. [19], this dependence falls outside the conditions typically used to justify standard resampling arguments (e.g., the nonparametric bootstrap), and can affect coverage in finite samples. Sample splitting or cross-validation can reduce this dependence by separating the construction of the data-based intervention rule from estimation of the target parameter, but these remedies introduce practical complexity: they can lead to split-specific versions of the estimand and complicate interpretation.
@@
Third, the performance of our approach depends on the stability of estimated intervention rule itself. In particular, the choice of the support level $\alpha$, the method for estimating conditional densities and the highest density regions. Instability in these components may introduce variability or sensitivity in finite samples, raising concerns about robustness and reproducibility across different applications.
@@
In light of these challenges, our work pursues a complementary goal through the feasible dose–response curve (FDRC): preserve the scientific objective of the original CDRC as closely as possible, while addressing positivity violations via an individualized feasible intervention rule. This yields a natural plug-in estimator that avoids extrapolation and provides valid inference under correct model specification, provided the intervention rule is reasonably stable. Although our approach does not deliver double robustness, it enhances interpretability and it improves interpretability and can reduce sensitivity to modeling choices by focusing estimation on well-supported regions. To evaluate the impact of estimand-data dependency on inference, we conducted a simulation study (Appendix G) comparing bootstrap coverage with and without sample splitting. The results suggest that, for the scenarios considered and in well-supported regions, bootstrap intervals can be reasonably accurate even without splitting.
@@
Finally, our method provides a principled approach for substituting infeasible interventions with their most feasible counterparts—thereby retaining an intervention-based interpretation closely related to the standard CDRC. Further research is needed to extend the framework to longitudinal settings and develop more robust techniques for estimating support regions, especially in high-dimensional and complex settings.
8 Software Implementation in the CICI Package
8.1 feasible(): Feasible-Intervention Estimation
The proposed feasible-intervention method is implemented in the R-CRAN package $CI$ [31, 32].
We illustrate usage with the package dataset EFV, which contains longitudinal treatment, outcome, and covariate measurements. For a one-time-point demonstration, we retain only baseline variables up to VL.0, yielding a one-time-point setup with treatment efv.0, outcome VL.0, and baseline covariates sex, metabolic, log_age, NRTI, and weight.0. In this reduced setting, there are no longitudinal covariate nodes (Lnodes = NULL).
We then estimate feasible interventions over target values abar_1tp = 0,1,…,10, summarize feasibility diagnostics, and visualize both the mean feasible mapping and non-overlap region.
        # A.X | One-time-point feasible intervention example (baseline only)              # Step 0: Load package and example data        # install.packages("CICI")        library(CICI)        data(EFV)              # Step 1: Prepare one-time-point data (baseline through VL.0)        # Variables kept: sex, metabolic, log_age, NRTI, weight.0, efv.0, VL.0        EFV_1tp <- EFV[, c("sex","metabolic","log_age","NRTI",        "weight.0","efv.0","VL.0")]              # Step 2: Define nodes for the one-time-point setting        Anodes <- "efv.0" # treatment at t = 0        Ynodes <- "VL.0" # outcome at t = 0 (illustrative baseline outcome)        Lnodes <- NULL # no longitudinal covariates in this reduced dataset              # Step 3: Define target intervention values        abar_1tp <- seq(0, 10, by = 1)              # Step 4: Estimate feasible interventions        m_1tp <- feasible(        X = EFV_1tp,        Anodes = Anodes,        Ynodes = Ynodes,        Lnodes = Lnodes,        abar = abar_1tp, # target treatment values to evaluate        alpha = 0.95, # feasibility level: retain top 95% density mass;        # the lowest ~5% tail is flagged as low-overlap        d.method = "parametric", # conditional density estimator g(a | L)        grid.size = 0.5, # internal grid spacing for density approximation        # smaller = finer resolution (more computation)        left.boundary = 0, # lower boundary of treatment support grid        right.boundary = 10 # upper boundary of treatment support grid        )              # Step 5: Inspect feasible values and overlap diagnostics        # Each m_1tp\$feasible[[k]] is an (n x 1) matrix:        # rows = individuals        # column = time point (here: efv.0 only)        # entries = individualized feasible value under strategy k (abar_1tp[k])              head(do.call(cbind, m_1tp\$feasible))        0 1 2 3 4 5 6 7 8 9 10        1 0.0 1.0 2 3 4 5 6 7 8 9 9.5        2 0.0 1.0 2 3 4 5 6 7 8 9 9.0        3 1.5 1.5 2 3 4 5 6 7 8 9 10.0        4 0.0 1.0 2 3 4 5 6 7 8 9 9.0        5 0.0 1.0 2 3 4 5 6 7 8 9 9.0        6 0.0 1.0 2 3 4 5 6 7 8 9 9.5              # Interpretation:        # After cbind, columns correspond to target strategies (abar_1tp = 0,...,10).        # Each cell is the individualized feasible value under that target.        # Values != the target (e.g., target 10 mapped to 9/9.5) indicate limited overlap.              # Strategy-level summaries        m_1tp # see ?print.feasible        summary(m_1tp) # see ?summary.feasible              # Step 6: Plot diagnostics        #        # Plot (a): which = "feasible"        # - x-axis: target intervention value a in abar_1tp        # - y-axis: mean feasible value E[d(a, .)] across individuals        #        # Plot (b): which = "nonoverlap"        # - low-density regions under fitted conditional treatment density g(a | L)              {internallinenumbers*} plot(m_1tp, which = "feasible") # mean feasible value across individuals, by target strategy        plot(m_1tp, which = "nonoverlap") # low-density/non-overlap region              # Step 7: Estimate and plot the feasible dose-response curve (FDRC)        #        # gformula() expects abar to represent the intervention values to plug in.        # Here we pass the individualized feasible interventions returned by feasible().              est_fdrc <- gformula(X = EFV_1tp,        Anodes = Anodes,        Ynodes = Ynodes,        Lnodes = NULL,        abar = m_1tp\$feasible # use results from Step 4)              plot(est_fdrc) # see ?plot.gformula
Figure 7a shows the mean feasible mapping, i.e., the average of individualized feasible values at each target intervention level. When the curve is close to the 45-degree identity line ($y=x$), target values are largely supported by the observed data. Systematic deviations from this identity indicate that targets are being remapped to nearby higher-density values, reflecting practical positivity constraints. Larger deviations imply stronger constraints.
Figure 7b shows treatment regions classified as low-density (non-overlap) under the fitted $g(a\mid H)$. These are areas with weak empirical support for direct intervention. Broader flagged regions indicate more severe overlap limitations, and flagged extremes typically correspond to where the mean feasible mapping departs most from $y=x$.
Figure 7c shows the estimated feasible dose–response curve (FDRC). The x-axis represents the target intervention value that we aim to set. At each target value, the FDRC reports the mean counterfactual outcome after assigning each individual to that target when it is realistically attainable for them, and otherwise to the closest attainable value returned by feasible(). Consequently, the curve typically agrees with the standard plug-in curve in well-attainable regions and departs from it primarily at extreme targets where many individuals are remapped to nearby values.
(a) Mean feasible intervention value (averaged over individuals) by target strategy.
(b) Estimated non-overlap (low-density) region under the fitted treatment density.
(c) Feasible dose–response curve (FDRC).
Figure 7: Software illustration for the one-time-point example. Panels (a)–(b) show feasibility diagnostics from feasible(): (a) the mean feasible mapping (average individualized feasible value at each target), and (b) intervention regions that are frequently unrealistic for some individuals (non-overlap ratio). Panel (c) shows the resulting feasible dose–response curve estimated via gformula() using the individualized feasible interventions; the x-axis denotes the target (attempted) intervention level, and outcomes are evaluated after mapping individuals who cannot realistically attain the target to their nearest attainable value.
8.2 Estimate the Conditional Density via Hazard-Binning
Consider discretizing the intervention $a$ of interest into $K$ bins, which may not necessarily be of the same width. Additionally, we can make the following observation: the probability of $a$ falling into a specific bin equals the probability of falling into the respective bin given that $a$ has not been placed into any of the previous bins times the probability of not falling into any of those those previous bins:
$P(a\in[\alpha_{k-1},\alpha_{k})|\mathbf{L})$ $=P(a\in[\alpha_{k-1},\alpha_{k})|A\geq\alpha_{k-1},\mathbf{L})\times$
$\ \ \prod_{j=1}^{k-1}\left\{1-P(a\in[\alpha_{j-1},\alpha_{j})|A\geq\alpha_{j-1},\mathbf{L})\right\}$ (14)
Under the assumption that within each interval $\alpha_{k}-\alpha_{k-1})$ the density is constant, we can estimate it as follows:
$f(a|\mathbf{l})=\frac{P(a\in[\alpha_{k-1},\alpha_{k})|\mathbf{L})}{\alpha_{k}-\alpha_{k-1}},\ \forall a\in[\alpha_{k}-\alpha_{k-1})$ (15)
One could estimate the probability of $a$ falling into the interval $\alpha_{k-1},\alpha_{k}),\ k=1,\cdots,K$, i.e. the left part of the above equation, using any regression or machine learning model. However, if density values are low in a specific interval this becomes very unstable. A safer option is to estimate the right hand side of (8.2) because the products of “one minus the conditional probabilities” impliy multiplication with “1” in low density intervals. To practically estimate the right hand side of (8.2), one has to re-format the data and fit a model “pooled over time” on it:
## Example of "pooled over bins" data for hazard-binning
## (three subjects, K = 3 bins on A)
id   A    L1    L2   k   alpha_start   alpha_end   in_bin
1  0.30  0.20  1.00  1      0.00         0.25        0
1  0.30  0.20  1.00  2      0.25         0.50        1
2  0.62 -0.50  0.30  1      0.00         0.25        0
2  0.62 -0.50  0.30  2      0.25         0.50        0
2  0.62 -0.50  0.30  3      0.50         0.75        1
3  0.10  1.10 -0.70  1      0.00         0.25        1
This data setup estimates the right-hand side of (8.2), specifically $P(a\in[\alpha_{k-1},\alpha_{k})\mid A\geq\alpha_{k-1},\mathbf{L})$ due to the implicit conditioning set: at each bin $k$, the pooled data restrict the sample to individuals with $A>\alpha_{k-1}$; thus, the regression model is fit on the corresponding risk set, and the estimated probabilities are conditional on “survival” through all previous bins. This mirrors discrete-time survival analyses, with $a$ playing the role of an event time and in_bin serving as the bin-specific failure indicator.
Note that with this reshaping the number of rows increases from $n$ to $\sum_{i=1}^{n}K_{i}$, where $K_{i}$ is the number of bins for which subject $i$ is still at risk (i.e. all $k$ with $\alpha_{k-1}\leq A_{i}$); in the worst case this is $nK$. In the implementation, we therefore report this pooled row count as an “effective sample size” when feasible() is called with verbose = TRUE and method = "hazardbinning".
The "hazardbinning" estimate the right-hand side of (8.2) through fitting pooled models on the reshaped data, then calculate the left-hand side of (8.2) and convert to density estimates (15).